A chemist needs 150 milliliters of a 43% solution but has only 13% and
58% solutions available. Find how many milliliters of each that should be
mixed to get the desired solution.
Can you show me how you formulated.. this is what I have so far.
43(150) = 13x + 58y

Answer by jorel1380(3719)

150x.43=64.5 mL
p+q=150
.13p+.58q=64.5
13p+58q=6450
13p+13q=1950
45q=4500
q=100
p=50
The chemist needs 50 mL of 13% solution and 100 mL of 58% solution..